OpenStudy (anonymous):

What is the simplified form of the expression? (m^-1 m^5 / m^-2)^-3 A. 3M^4/M^-2 B. 1/M^18 C. M^18 D. -M^216

OpenStudy (lgbasallote):

\[\Large (\frac{m^{-1} m^5}{m^{-2}})^{-3} \implies (\frac{m^{-1+5}}{m^{-2}})^{-3}\] \[\Large \implies (\frac{m^4}{m^{-2}})^{-3} \implies (m^{4 - (-2)})^{-3} \implies (m^{4 +2})^{-3} \] \[\Large \implies (m^6)^{-3} \implies m^{-18} \implies \frac{1}{m^{18}}\] finally one of your options match p

OpenStudy (lgbasallote):

i hope you udnerstood my solution though

OpenStudy (anonymous):

yes i did thanks